# Hengwei Bian1,2,    Lingdong Kong1,3   Haozhe Xie4   Liang Pan1,†,‡   Yu Qiao1   Ziwei Liu4

### v2

@@ A.3 Model Details
Discussion on VAE Structure Improvements. Some prior work utilizes sparse 3D structures to enhance the efficiency of their backbones. For example, XCube (Ren et al., 2024b) employs a fully sparse 3D encoder, significantly improving model efficiency. Similarly, our VAE could potentially improve the 3D convolutional feature extractor $\bm{f}_{\theta}(\cdot)$ by adopting sparse convolution. However, using sparse convolution offers only limited efficiency gains, as convolution accounts for only a small portion of our VAE. Moreover, like XCube, we cannot apply sparse convolution in our decoder. In the future, we plan to explore more efficient operations to further optimize our 3D backbone.
A.4 Classifier-Free Guidance
Classifier-Free Guidance (CFG) (Ho & Salimans, 2022) could improve the performance of conditional generative models without relying on an external classifier. Specifically, during training, the model simultaneously learns both conditional generation $p(x|c)$ and unconditional generation $p(x)$, and guidance during sampling is provided by the following equation:
$\hat{x}_{t}=(1+w)\cdot\hat{x}_{t}(c)-w\cdot\hat{x}_{t}(\emptyset)~{},$ (12)
@@ A.5 Downstream Applications
Layout. To control object placement in the scene, we train a model capable of generating vehicle dynamics based on a bird’s-eye view sketch. We apply semantic filtering to the bird’s-eye view of the input scene, marking regions with vehicles as $1$ and regions without vehicles as 00. Pooling this binary image provides layout information as a $T\times H\times W$ tensor from the bird’s-eye perspective. The layout is padded to match the size of the HexPlane, ensuring that the positional encoding of the bird’s-eye layout aligns with the $XY$ plane. DiT learns the correspondence between the layout and vehicle semantics using the same conditional injection method applied to the HexPlane.
@@
Inpaint. We demonstrate that our model can handle versatile applications by training a conditional DiT for the previous tasks. Extending our exploration of downstream applications, and inspired by (Lee et al., 2024), we leverage the 2D structure of our latent space and the explicit modeling of each dimension to highlight our model’s ability to perform inpainting on 4D scenes. During DiT sampling, we define a 2D mask $m\in\mathbb{R}^{X\times Y}$ on the $XY$ plane, which is extended across all dimensions to mask specific regions of the HexPlane.
@@
$\mathcal{H}_{t}=m\odot\mathcal{H}_{t}+(1-m)\odot\mathcal{H}_{t}^{\text{in}}~{},$ (13)
@@
Appendix B B Additional Quantitative Results
@@
B.1 Per-Class Generation Results
We include the class-wise IoU scores of OccSora (Wang et al., 2024) and our proposed DynamicCity framework on CarlaSC (Wilson et al., 2022). As shown in Tab. 7, our results demonstrate higher IoU across all classes, indicating that our VAE reconstruction achieves minimal information loss. Additionally, our model does not exhibit significantly low IoU for any specific class, proving its ability to effectively handle class imbalance.
Table 7: Comparisons of Per-Class IoU Scores. We compared the performance of OccSora (Wang et al., 2024), and our DynamicCity framework on CarlaSC (Wilson et al., 2022) across $10$ semantic classes. The scene resolution is $128\times 128\times 8$. The sequence lengths are $4$, $8$, $16$, and $32$, respectively.
@@ B.1 Per-Class Generation Results
Resolution: $128\times 128\times 8$    Sequence Length: $4$
@@
Resolution: $128\times 128\times 8$    Sequence Length: $8$
@@
B.2 Occupancy Forecasting Results
We train our HexPlane conditional generation pipeline on Occ3D-nuScenes (Tian et al., 2023) as an occupancy forecasting model. We set $T=4$ to ensure the model receives a HexPlane with a context length of 2 seconds, aligning with OccWorld (Zheng et al., 2024a), and generates the next 2 seconds for evaluation. As shown in Tab. 8, our model outperforms OccWorld on most metrics.
Table 8: 4D Occupancy Forecasting Performance. We compare the performance of OccWorld (Zheng et al., 2024a) and our proposed DynamicCity framework on Occ3D-nuScenes (Tian et al., 2023).
@@ B.2 Occupancy Forecasting Results
B.3 User Study
We conduct a user study comparing OccSora (Wang et al., 2024) with our proposed DynamicCity. The study includes 20 samples, with 10 from each method. Participants rate each sample on four metrics: 1) overall quality, 2) time consistency, 3) background quality, and 4) foreground quality. Ratings range from 1 to 5, with 5 being the highest. We collect results from 42 volunteers and get 840 valid scores in total, as shown in Tab. 9. Our method receives better user feedback across all metrics.
Table 9: User Study Results. We conduct user study comparing OccSora (Wang et al., 2024) and DynamicCity. The rating is of scale 1-5, the higher the better.
@@ B.3 User Study
B.4 Model Stats
We compare the training speed, inference speed, training VRAM, and inference VRAM of OccSora (Wang et al., 2024) and DynamicCity. The results are presented in Tab. 10, Tab. 11, Tab. 12, and Tab. 13. While some of our models may be slightly slower and consume more memory compared to OccSora, they achieve significantly better performance. We also compare the total model size of OccSora and our model in Tab. 14. Our model is significantly smaller than OccSora while achieving superior performance.
Table 10: VAE Model Statistics on CarlaSC Dataset (Wilson et al., 2022). We compare the training time, inference time, training VRAM, inference VRAM of OccSora (Wang et al., 2024) and our DynamicCity.
Method Training Time (s)$\downarrow$ Inference Time (s)$\downarrow$ Training VRAM (G)$\downarrow$ Inference VRAM (G)$\downarrow$
@@ B.4 Model Stats
Table 11: VAE Model Statistics on Occ3D-Waymo Dataset (Tian et al., 2023). We compare the training time, inference time, training VRAM, inference VRAM of OccSora (Wang et al., 2024) and our DynamicCity.
Method Training Time (s)$\downarrow$ Inference Time (s)$\downarrow$ Training VRAM (G)$\downarrow$ Inference VRAM (G)$\downarrow$
@@
Table 12: DiT Model Statistics on CarlaSC Dataset (Wilson et al., 2022). We compare the training time, inference time, training VRAM, inference VRAM of OccSora (Wang et al., 2024) and our DynamicCity.
Method Training Time (s)$\downarrow$ Inference Time (s)$\downarrow$ Training VRAM (G)$\downarrow$ Inference VRAM (G)$\downarrow$
@@
Table 13: DiT Model Statistics on Occ3D-Waymo Dataset (Tian et al., 2023). We compare the training time, inference time, training VRAM, inference VRAM of OccSora (Wang et al., 2024) and our DynamicCity.
Method Training Time (s)$\downarrow$ Inference Time (s)$\downarrow$ Training VRAM (G)$\downarrow$ Inference VRAM (G)$\downarrow$
@@
Table 14: Model Size. We compare the total model size of OccSora (Wang et al., 2024) and our DynamicCity.
@@
B.5 Comparisons with SemCity
We compare the generation quality of SemCity (Lee et al., 2024) and our DynamicCity in Tab. 15. Despite using a more compact latent representation and generating dynamic scenes, our model outperforms SemCity on most metrics.
Table 15: Comparisons of 2D and 3D Evaluation Metrics. We report the Inception Score (IS), Fréchet Inception Distance (FID), Kernel Inception Distance (KID), and the Precision (P) and Recall (R) rates for SemCity (Lee et al., 2024) and our method in both the 2D and 3D spaces.
[…]
@@ B.5 Comparisons with SemCity
Appendix C C Additional Qualitative Results
In this section, we provide additional qualitative results on the Occ3D-Waymo (Tian et al., 2023) and CarlaSC (Wilson et al., 2022) datasets to demonstrate the effectiveness of our approach.
C.1 Unconditional Dynamic Scene Generation
@@
Figure 9: Unconditional Dynamic Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on the Occ3D-Waymo (Tian et al., 2023) dataset. Best viewed in colors and zoomed-in for additional details.
Figure 10: Unconditional Dynamic Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on the CarlaSC (Wilson et al., 2022) dataset. Best viewed in colors and zoomed-in for additional details.
C.2 HexPlane-Guided Generation
@@
Figure 11: HexPlane-Guided Generation Results. We provide qualitative examples of a total of $64$ consectutive frames generated by DynamicCity on the Occ3D-Waymo (Tian et al., 2023) dataset. Best viewed in colors and zoomed-in for additional details.
C.3 Layout-Guided Generation
@@
Figure 12: Layout-Guided Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on the Occ3D-Waymo (Tian et al., 2023) dataset. Best viewed in colors and zoomed-in for additional details.
C.4 Command- & Trajectory-Guided Generation
@@
Figure 13: Command-Guided Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated under the command RIGHT by DynamicCity on the CarlaSC (Wilson et al., 2022) dataset. Best viewed in colors and zoomed-in for additional details.
Figure 14: Trajectory-Guided Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on the CarlaSC (Wilson et al., 2022) dataset. Best viewed in colors and zoomed-in for additional details.
C.5 Dynamic Inpainting
@@
Figure 15: Dynamic Inpainting Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on the CarlaSC (Wilson et al., 2022) dataset. Best viewed in colors and zoomed-in for additional details.
C.6 Comparisons with OccSora
We compare our qualitative results with OccSora (Wang et al., 2024) in Fig. 16, using a similar scene. It is evident that our result presents a realistic dynamic scene, with straight roads and complete objects and environments. In contrast, OccSora’s result displays unreasonable semantics, such as a pedestrian in the middle of the road, broken vehicles, and a lack of dynamic elements. This comparison highlights the effectiveness of our method.
Figure 16: Comparisons of Dynamic Scene Generation. We provide qualitative examples of a total of $16$ consectutive frames generated by OccSora (Wang et al., 2024) and our proposed DynamicCity framework on the CarlaSC (Wilson et al., 2022) dataset. Best viewed in colors and zoomed-in for additional details.
C.7 Dynamic Outpainting
@@
Figure 17: Dynamic Outpainting Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on the CarlaSC (Wilson et al., 2022) dataset. Best viewed in colors and zoomed-in for additional details.
C.8 Single Frame Occupancy Conditional Generation
@@
Figure 18: Single Frame Occupancy Conditional Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on the CarlaSC (Wilson et al., 2022) dataset. Best viewed in colors and zoomed-in for additional details.
Appendix D D Potential Societal Impact & Limitations
@@
D.1 Societal Impact
Our approach’s ability to generate high-quality 4D occupancy holds the potential to significantly impact various domains, particularly autonomous driving, robotics, urban planning, and smart city development. By creating realistic, large-scale dynamic scenes, our model can aid in developing more robust and safe autonomous systems. These systems can be better trained and evaluated against diverse scenarios, including rare but critical edge cases like unexpected pedestrian movements or complex traffic patterns, which are difficult to capture in real-world datasets. This contribution can lead to safer autonomous vehicles, reducing traffic accidents, and improving traffic efficiency, ultimately benefiting society by enhancing transportation systems.
@@ D.2 Broader Impact
However, it is crucial to emphasize that synthetic data should be used responsibly. As our model generates highly realistic scenes, there is a risk that reliance on synthetic data could lead to models that fail to generalize effectively in real-world settings, especially if the generated scenes do not capture the full diversity or rare conditions found in real environments. Hence, it’s important to complement synthetic data with real-world data and ensure transparency when using synthetic data in model training and evaluation.
D.3 Known Limitations
@@
Appendix E E Public Resources Used
@@
E.1 Public Datasets Used
E.2 Public Implementations Used

### v3

@@ 7.3 Model Details
Discussion on VAE Structure Improvements. Some prior work utilizes sparse 3D structures to enhance the efficiency of their backbones. For example, XCube [30] employs a fully sparse 3D encoder, significantly improving model efficiency. Similarly, our VAE could potentially improve the 3D convolutional feature extractor $\bm{f}_{\theta}(\cdot)$ by adopting sparse convolution. However, using sparse convolution offers only limited efficiency gains, as convolution accounts for only a small portion of our VAE. Moreover, like XCube, we cannot apply sparse convolution in our decoder. In the future, we plan to explore more efficient operations to further optimize our 3D backbone.
7.4 Classifier-Free Guidance
Classifier-Free Guidance (CFG) [12] could improve the performance of conditional generative models without relying on an external classifier. Specifically, during training, the model simultaneously learns both conditional generation $p(x|c)$ and unconditional generation $p(x)$, and guidance during sampling is provided by the following equation:
$\hat{x}_{t}=(1+w)\cdot\hat{x}_{t}(c)-w\cdot\hat{x}_{t}(\emptyset)\penalty 10000\ ,$ (12)
@@ 7.5 Downstream Applications
Layout. To control object placement in the scene, we train a model capable of generating vehicle dynamics based on a bird’s-eye view sketch. We apply semantic filtering to the bird’s-eye view of the input scene, marking regions with vehicles as $1$ and regions without vehicles as 0. Pooling this binary image provides layout information as a $T\times H\times W$ tensor from the bird’s-eye perspective. The layout is padded to match the size of the HexPlane, ensuring that the positional encoding of the bird’s-eye layout aligns with the $XY$ plane. DiT learns the correspondence between the layout and vehicle semantics using the same conditional injection method applied to the HexPlane.
@@
Inpaint. We demonstrate that our model can handle versatile applications by training a conditional DiT for the previous tasks. Extending our exploration of downstream applications, and inspired by [17], we leverage the 2D structure of our latent space and the explicit modeling of each dimension to highlight our model’s ability to perform inpainting on 4D scenes. During DiT sampling, we define a 2D mask $m\in\mathbb{R}^{X\times Y}$ on the $XY$ plane, which is extended across all dimensions to mask specific regions of the HexPlane.
@@
$\mathcal{H}_{t}=m\odot\mathcal{H}_{t}+(1-m)\odot\mathcal{H}_{t}^{\text{in}}\penalty 10000\ ,$ (13)
@@
8 Additional Quantitative Results
@@
8.1 Per-Class Generation Results
We include the class-wise IoU scores of OccSora [41] and our proposed DynamicCity framework on CarlaSC [42]. As shown in Tab. 7, our results demonstrate higher IoU across all classes, indicating that our VAE reconstruction achieves minimal information loss. Additionally, our model does not exhibit significantly low IoU for any specific class, proving its ability to effectively handle class imbalance.
Table 7: Comparisons of Per-Class IoU Scores. We compared the performance of OccSora [41], and our DynamicCity framework on CarlaSC [42] across $10$ semantic classes. The scene resolution is $128\times 128\times 8$. The sequence lengths are 44, 88, $16$, and $32$, respectively.
@@ 8.1 Per-Class Generation Results
Resolution: $128\times 128\times 8$ Sequence Length: 44
@@
Resolution: $128\times 128\times 8$ Sequence Length: 88
@@
8.2 Occupancy Forecasting Results
We train our HexPlane conditional generation pipeline on Occ3D-nuScenes [40] as an occupancy forecasting model. We set $T=4$ to ensure the model receives a HexPlane with a context length of 2 seconds, aligning with OccWorld [48], and generates the next 2 seconds for evaluation. As shown in Tab. 8, our model outperforms OccWorld on most metrics.
Table 8: 4D Occupancy Forecasting Performance. We compare the performance of OccWorld [48] and our proposed DynamicCity framework on Occ3D-nuScenes [40].
@@ 8.2 Occupancy Forecasting Results
8.3 User Study
We conduct a user study comparing OccSora [41] with our proposed DynamicCity. The study includes 20 samples, with 10 from each method. Participants rate each sample on four metrics: 1) overall quality, 2) time consistency, 3) background quality, and 4) foreground quality. Ratings range from 1 to 5, with 5 being the highest. We collect results from 42 volunteers and get 840 valid scores in total, as shown in Tab. 9. Our method receives better user feedback across all metrics.
Table 9: User Study Results. We conduct user study comparing OccSora [41] and DynamicCity. The rating is of scale 1-5, the higher the better.
@@ 8.3 User Study
8.4 Model Stats
We compare the training speed, inference speed, training VRAM, and inference VRAM of OccSora [41] and DynamicCity. The results are presented in Tab. 10, Tab. 11, Tab. 12, and Tab. 13. While some of our models may be slightly slower and consume more memory compared to OccSora, they achieve significantly better performance. We also compare the total model size of OccSora and our model in Tab. 14. Our model is significantly smaller than OccSora while achieving superior performance.
Table 10: VAE Model Statistics on CarlaSC Dataset [42]. We compare the training time, inference time, training VRAM, inference VRAM of OccSora [41] and our DynamicCity.
[…]
@@ 8.4 Model Stats
Table 11: VAE Model Statistics on Occ3D-Waymo Dataset [40]. We compare the training time, inference time, training VRAM, inference VRAM of OccSora [41] and our DynamicCity.
[…]
@@
Table 12: DiT Model Statistics on CarlaSC Dataset [42]. We compare the training time, inference time, training VRAM, inference VRAM of OccSora [41] and our DynamicCity.
[…]
@@
Table 13: DiT Model Statistics on Occ3D-Waymo Dataset [40]. We compare the training time, inference time, training VRAM, inference VRAM of OccSora [41] and our DynamicCity.
[…]
@@
Table 14: Model Size. We compare the total model size of OccSora [41] and our DynamicCity.
@@
8.5 Comparisons with SemCity
We compare the generation quality of SemCity [17] and our DynamicCity in Tab. 15. Despite using a more compact latent representation and generating dynamic scenes, our model outperforms SemCity on most metrics.
Table 15: Comparisons of 2D and 3D Evaluation Metrics. We report the Inception Score (IS), Fréchet Inception Distance (FID), Kernel Inception Distance (KID), and the Precision (P) and Recall (R) rates for SemCity [17] and our method in both the 2D and 3D spaces.
Method $^{{\color[rgb]{0.890625,0.421875,0.76171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.890625,0.421875,0.76171875}\textbf{2D}}} Metric$$^{{\color[rgb]{0.5,0.625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.625,0.8828125}\textbf{3D}}}$
@@ 8.5 Comparisons with SemCity
9 Additional Qualitative Results
In this section, we provide additional qualitative results on the Occ3D-Waymo [40] and CarlaSC [42] datasets to demonstrate the effectiveness of our approach.
9.1 Unconditional Dynamic Scene Generation
@@
Figure 9: Unconditional Dynamic Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on Occ3D-Waymo [40]. Best viewed in colors and zoomed-in for additional details.
Figure 10: Unconditional Dynamic Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on CarlaSC [42]. Best viewed in colors and zoomed-in for additional details.
9.2 HexPlane-Guided Generation
@@
Figure 11: HexPlane-Guided Generation Results. We provide qualitative examples of a total of $64$ consectutive frames generated by DynamicCity on Occ3D-Waymo [40]. Best viewed in colors and zoomed-in for additional details.
9.3 Layout-Guided Generation
@@
Figure 12: Layout-Guided Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on Occ3D-Waymo [40]. Best viewed in colors and zoomed-in for additional details.
9.4 Command- & Trajectory-Guided Generation
@@
Figure 13: Command-Guided Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated under the command RIGHT by DynamicCity on CarlaSC [42]. Best viewed in colors and zoomed-in for additional details.
Figure 14: Trajectory-Guided Scene Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on CarlaSC [42]. Best viewed in colors and zoomed-in for additional details.
9.5 Dynamic Inpainting
@@
Figure 15: Dynamic Inpainting Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on CarlaSC [42]. Best viewed in colors and zoomed-in for additional details.
9.6 Comparisons with OccSora
We compare our qualitative results with OccSora [41] in Fig. 16, using a similar scene. It is evident that our result presents a realistic dynamic scene, with straight roads and complete objects and environments. In contrast, OccSora’s result displays unreasonable semantics, such as a pedestrian in the middle of the road, broken vehicles, and a lack of dynamic elements. This comparison highlights the effectiveness of our method.
Figure 16: Comparisons of Dynamic Scene Generation. We provide qualitative examples of a total of $16$ consectutive frames generated by OccSora [41] and our proposed DynamicCity framework on CarlaSC [42]. Best viewed in colors and zoomed-in for additional details.
9.7 Dynamic Outpainting
@@
Figure 17: Dynamic Outpainting Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on CarlaSC [42]. Best viewed in colors and zoomed-in for additional details.
9.8 Single Frame Occupancy Conditional Generation
@@
Figure 18: Single Frame Occupancy Conditional Generation Results. We provide qualitative examples of a total of $16$ consectutive frames generated by DynamicCity on CarlaSC [42]. Best viewed in colors and zoomed-in for additional details.
10 Potential Societal Impact & Limitations
@@
10.1 Societal Impact
Our approach’s ability to generate high-quality 4D occupancy holds the potential to significantly impact various domains, particularly autonomous driving, robotics, urban planning, and smart city development. By creating realistic, large-scale, dynamic scenes, our model can aid in developing more robust and safe autonomous systems. These systems can be better trained and evaluated against diverse scenarios, including rare but critical edge cases like unexpected pedestrian movements or complex traffic patterns, which are difficult to capture in real-world datasets. This contribution can lead to safer autonomous vehicles, reducing traffic accidents, and improving traffic efficiency, ultimately benefiting society by enhancing transportation systems.
@@ 10.2 Broader Impact
However, it is crucial to emphasize that synthetic data should be used responsibly. As our model generates highly realistic scenes, there is a risk that reliance on synthetic data could lead to models that fail to generalize effectively in real-world settings, especially if the generated scenes do not capture the full diversity or rare conditions found in real environments. Hence, it is important to complement synthetic data with real-world data and ensure transparency when using synthetic data in model training and evaluation.
10.3 Known Limitations
@@
11 Public Resources Used
@@
11.1 Public Datasets Used
  • nuScenes [^1] https://www.nuscenes.org/nuscenes ........................................................................................................................................................................CC BY-NC-SA 4.0
  • nuScenes-devkit [^2] https://github.com/nutonomy/nuscenes-devkit ........................................................................................................................................................................Apache License 2.0
  • Waymo Open Dataset [^3] https://waymo.com/open ........................................................................................................................................................................Waymo Dataset License
  • CarlaSC [^4] https://umich-curly.github.io/CarlaSC.github.io. ........................................................................................................................................................................MIT License
  • Occ3D [^5] https://tsinghua-mars-lab.github.io/Occ3D. ........................................................................................................................................................................MIT License
11.2 Public Implementations Used
  • SemCity [^6] https://github.com/zoomin-lee/SemCity. ........................................................................................................................................................................Unknown
  • OccSora [^7] https://github.com/wzzheng/OccSora. ........................................................................................................................................................................Apache License 2.0
  • MinkowskiEngine [^8] https://github.com/NVIDIA/MinkowskiEngine. ........................................................................................................................................................................MIT License
  • TorchSparse [^9] https://github.com/mit-han-lab/torchsparse. ........................................................................................................................................................................MIT License
  • SPVNAS [^10] https://github.com/mit-han-lab/spvnas. ........................................................................................................................................................................MIT License
  • spconv [^11] https://github.com/traveller59/spconv. ........................................................................................................................................................................Apache License 2.0